5. An epidemiological study looked at the relationship between earbud use and hearing loss. Data for number of earbud users appear in the table below. Test for an association betweern earbud use and hearing loss at the 5% level of significance. (4 points) Daily Earbud Use (min.) Hearing Loss> 60 60-120 120-180 >180 None Slight Moderate 562 89 698 90 26 239 29 201
Answer #1

here null hypothesis:Ho: earbud use time and hearing loss are not related

alternate hypothesis:Ha:  earbud use time and hearing loss are related

degree of freedom =(row-1)*(column-1)=(3-1)*(4-1)=6

for 6 df and 0.05 level of significance ; critical value \chi ^{2} =12.592

applying chi square test of independence:

Observed OBS <60 60-120 120-180 >180 Total
none 27 19 6 6 58
slight 35 27 10 10 82
moderate 26 18 10 10 64
Total 88 64 26 26 204
Expected Ei=?row*?column/?total <60 60-120 120-180 >180 Total
none 25.020 18.196 7.392 7.392 58
slight 35.373 25.725 10.451 10.451 82
moderate 27.608 20.078 8.157 8.157 64
Total 88 64 26 26 204
chi square =(Oi-Ei)2/Ei <60 60-120 120-180 >180 Total
none 0.16 0.04 0.26 0.26 0.72
slight 0.00 0.063 0.019 0.019 0.11
moderate 0.09 0.215 0.416 0.416 1.14
Total 0.25 0.31 0.70 0.70 1.964

for test statistic 1.964 is less than crtiical value; we fail to reject null hypothesis

we do not have evidence to conclude that earbud use and hearing loss are related,